nikkixmarie:

If TU = 3x-12 and SU = 4x + 2, find the value of x. Then find the length of TU and SU. If ST= x + 3 and TU= 4x-6, find the value of x. Then find the length of ST and TU.

nikkixmarie:

@TheSmartOne

TheSmartOne:

There must be a diagram that came with the question

nikkixmarie:

there isnt at all

TheSmartOne:

Any other information? Is TU and SU the same length

nikkixmarie:

i dont know thats all it asked me

TheSmartOne:

There has to be like one sentence before the question at least

nikkixmarie:

ohhh ill take a pic my bad there is a diagram

nikkixmarie:

give me a sec

nikkixmarie:

1 attachment
TheSmartOne:

So ST = TU because they're bot the same length SU is the whole line and if you divide it by two you get the length of ST and TU

TheSmartOne:

so SU/2 = TU so plug in the expressions and solve for x (4x+2)/2 = 3x-12

TheSmartOne:

And for question #3 you just set ST = TU and solve for x And once you get the value of x, you plug it in to the expressions and get the value. And you should get the same number for ST and TU

nikkixmarie:

x = 13

nikkixmarie:

and then x=6

nikkixmarie:

how do i plug it in the expression?

TheSmartOne:

First one is correct x=6 is wrong #3) x+3 = 4x-6 solve for x

TheSmartOne:

you plug it in just like you did for question #1, I'm not sure how you correctly found out x=6 for that question but you plugged it in correctly :)

nikkixmarie:

so thats how i would get ST and TU

nikkixmarie:

is it 39?

nikkixmarie:

@Ultrilliam

nikkixmarie:

can i plz get help

nikkixmarie:

@Vocaloid

TheSmartOne:

You plug in x=13 for question #2 so essentially SU = 4(13)+2 = ?? TU = 3(13)- 12 = ??